# Circuits and Backdoors: Five Shades of the SETH

### v1

@@ 5 Modulators to Logarithmic Pathwidth and 2-SAT Backdoors
[…]
@@
[…]
@@
[…]
Proof 5.3 (Proof of Theorem 5.1).
We split the proof into three parts. In Lemma 5.4 we show an equivalence between 2-SAT backdoors and Ann-s-t-Non-Reach; in Lemma 5.10 we show an equivalence between SAT with a modulator to logarithmic pathwidth and Ann-s-t-Reach; and in Lemma 5.12 we show that Ann-s-t-Reach and Ann-s-t-Non-Reach have the same complexity by revisiting the proof of the Immerman–Szelepcsényi theorem.
@@
Proof 5.5.
The proof is broken down into Lemma 5.6 and Lemma 5.8.
@@
We use a standard reduction of non-reachability to 2-SAT, except we have to take into account the arc annotations. Suppose we are given a Ann-s-t-Non-Reach instance. Let $G=(V,E)$ be the given DAG on $n$ vertices. We construct a CNF formula with variables the variables of $M$, given with the Ann-s-t-Non-Reach instance, to which we add a variable $x_{v}$ for each $v\in V$. We will set $M$ to be the backdoor set. Now, for each arc $(v_{1},v_{2})\in E$ we add the clause ($(\neg x_{v_{1}}\lor x_{v_{2}})$ to the formula if the arc is not annotated, otherwise we add the clause ($(\neg x_{v_{1}}\lor x_{v_{2}}\lor\neg\ell)$ if the arc is annotated with the literal $\ell$. We also add the unit clauses $(x_{s})$ and ($(\neg x_{t})$. This completes the construction.
We now claim that if the formula we constructed is satisfiable, using the same assignment to $M$ yields a digraph where there is no path from $s$ to $t$. Indeed, suppose that for the same assignment there is a path from $s$ to $t$, meaning that the assignment sets to True all the literals used as annotations to the edges of the path. This implies that in the corresponding closes, the backdoor literals we have added are set to False by the satisfying assignment. But then this would create a chain of implications given $x_{s}\to x_{t}$, which in conjunction with the two unit clauses would falsify the formula.
@@
Suppose now that the original CNF is satisfiable. We want to show that by using the same assignment to $M$ we eliminate all $s\to t$ paths. Indeed, suppose we find such a path, which for some $i\in[n]$ must go through $D_{i,0},D_{i,1}$. In particular, this means we have a path from $x_{i}$ to $\neg x_{i}$ and a path from $\neg x_{i}$ to $x_{i}$ in $D$ while respecting the annotations. However, recall that for each clause we have added a path of annotated arcs containing the negations of all backdoor literals of the clause, meaning that if we can follow an arc, the assignment to $M$ must be setting all the backdoor literals to False. We conclude that the assignment cannot be satisfying to the formula, as we can infer the implications $(x_{i}\to\neg x_{i})$ and ($(\neg x_{i}\to x_{i})$.
@@ 5.2 Log-Pathwidth and Reachability
[…]
@@
Let us now move to the other half of the lemma. Suppose there is an algorithm deciding satisfiability with a modulator $M$ to pathwidth $O(\log|\phi|)$ and we want to use this algorithm to solve Ann-s-t-Reach. We construct a formula which retains the set $M$ of annotation variables as the modulator. Given a digraph $G=(V,E)$ with $n$ vertices we assume without loss of generality that $n$ is a power of $2$ (otherwise add some dummy vertices). We construct a formula with $n\log n$ extra variables, called $x_{i,j}$, $i\in[n],j\in[\log n]$. The intended meaning is that the $\log n$ variables $x_{i,1},\ldots,x_{i,\log n}$ encode the index of the $i$-th vertex we would visit in an $s\to t$ path. To enforce this meaning we assume the vertices of $G$ are numbered from 00 to $n-1$ and with each vertex we associate a distinct truth assignment to $\log n$ variables by reading these variables as a binary number. We add clauses enforcing the variables $x_{1,j},j\in[\log n]$ to take the value corresponding to $s$ and the variables $x_{n,j},j\in[\log n]$ to take the value corresponding to $t$. For each $i\in[n-1]$ we add at most $n^{2}$ clauses to the formula as follows: for each $(u,v)\not\in E$ such that $u\neq v$, we add a clause ensuring that if the $x_{i,j}$ variables encode $u$ and the $x_{i+1,j}$ variables encode $v$, then the formula is not satisfied; for each $(u,v)\in E$ which is annotated with a literal $\ell$ we add the same clause (which is falsified if $x_{i,j},x_{i+1,j}$ encode $u,v$), but we add to this clause the literal $\ell$. If $(u,v)\in E$ is not annotated or if $u=v$, we do not add a clause. It is not hard to see that, if we remove the variables of $M$, this formula has pathwidth $O(\log n)$, because we can make $n-1$ bags, each containing $x_{i,j},x_{i+1,j},j\in[\log n]$ and this covers all edges.
To argue for correctness, if there is an $s\to t$ path in the original instance for some assignment to $M$ we use the same assignment in $\phi$ and extend it to a satisfying assignment as follows: if the path visits $r\leq n$ vertices in total, we set for each $i\in[r]$ the assignment to $x_{i,j},j\in[\log n]$ to encode the $i$-th vertex of the path; all remaining groups encode $t$. This satisfies the formula. For the converse direction, we can infer from a satisfying assignment to the formula a path of length at most $n-1$ by interpreting the assignment to $x_{i,j},j\in[\log n]$ as the encoding of the $i$-th vertex of the path, and the clauses we have added ensure that the path is valid, that is we are only followed to stay in the same vertex, follow an arc without annotation, or follow an arc whose annotation is set to True by the assignment to $M$.
@@ 5.3 Reachability and Non-Reachability
Before we begin, let us slightly edit $G_{1}$. Take $n=|V_{1}|$ copies of $V_{1}$, constructing for each $v\in V_{1}$ the vertices $v^{1},\ldots,v^{n}$. For each $i\in[n-1]$ and $v\in V_{1}$ we add the arc $(v^{i},v^{i+1})$ without annotation; for each arc $(u,v)$ and each $i\in[n-1]$ we add the arc $(u^{i},v^{i+1})$, maintaining the annotation of $(u,v)$. Finally, for the vertex $s^{n-1}$ and all $v\in V_{i}\setminus\{t\}$ we add the non-annotated arc $(s^{n-1},v^{n})$. This transformation has achieved the following: for each assignment to $M$, an $s\to t$ path exists in $G_{1}$ if and only if an $s^{1}\to t^{n}$ path exists in the new graph; there is a path from $s_{1}$ to $v^{n}$ for all $v\in V_{1}\setminus\{t\}$. As a result the question of whether an $s\to t$ path exists in $G_{1}$ is equivalent to asking whether the number of vertices reachable from $s^{1}$ in $\{v^{n}\ |\ v\in V_{1}\}$ is equal to $n$. Call the new graph $G_{1}^{\prime}$ and we want to construct $G_{2}$ such that for all assignments to $M$, there is a path $s_{2}\to t_{2}$ in $G_{2}$ if and only if $n-1$ vertices of $\{v^{n}\ |\ v\in V_{1}\}$ are reachable from $s^{1}$ in $G_{1}^{\prime}$. In the remainder, when we refer to the $i$-th layer of $G_{1}^{\prime}$ we will mean the set $\{v^{i}\ |v\in V_{1}\}$. Observe that $G_{1}^{\prime}$ has the property that all arcs go from a layer $i$ to a layer $i+1$.
We begin the construction of $G_{2}$ with $n^{2}$ vertices $s_{i,j}$ for $i,j\in[n]$, plus a source $s_{2}$ and a sink $t_{2}$. The intended meaning is that (for all assignments to $M$) $s_{i,j}$ will be reachable from $s_{2}$ if and only if there are exactly $j$ vertices of $G_{1}^{\prime}$ in layer $i$ which are reachable from $s^{1}$. We add a non-annotated arc $(s_{n,n-1},t_{2})$ and a non-annotated arc $(s_{2},s_{1,1})$. For each $i,j\in[n]$ we want to construct a checker gadget $C_{i,j}$ such that there will be an arc from $s_{i,j}$ to $C_{i,j}$ and arcs from $C_{i,j}$ to $s_{i+1,j^{\prime}}$ for some $j^{\prime}\in[n]$. We want to achieve the following property for all assignments $\sigma$ to $M$. Suppose that under some assignment $\sigma$, there are exactly $j$ vertices of layer $i$ of $G_{1}^{\prime}$ reachable from $s^{1}$. Then, there will be a path from $s_{i,j}$ to $s_{i+1,j^{\prime}}$ through $C_{i,j}$ if and only if the number of vertices of layer $i+1$ of $G_{1}^{\prime}$ reachable from $s^{1}$ is $j^{\prime}$. If we achieve this property we are done, as any path starts from $s_{1,1}$ and it is the case that exactly $1$ vertex of layer $1$ is reachable from $s^{1}$ ($s^{1}$ itself), so we can proceed by induction and we see that there is a path to $s_{n,n-1}$ (hence to $t_{2}$) if and only if $n-1$ vertices of the last layer of $G_{1}^{\prime}$ are reachable. Hence, there will be a $s_{2}\to t_{2}$ path in $G_{2}$ if and only if there is no $s_{1}\to t_{1}$ path in $G_{1}$.
@@ 6 Circuit-SAT and Horn Backdoors
Before we go on let us give some intuition. Similarly to Section 5, it will be helpful to recall some basic facts from computational complexity theory. For the first two statements, we are dealing with the satisfiability problem for circuits and SAT with a Horn backdoor. Intuitively, it is natural to expect these two problems to be equivalent because once we fix either the input to a given circuit or the assignment to the Horn backdoor, deciding if the circuit is satisfied or if the backdoor assignment can be extended to a satisfying assignment are problems with the same classical complexity, that is, they are P-complete [39]. As in Section 5 this intuition is not in itself sufficient to imply the equivalence, because what we want is not to claim that once we have an assignment, checking it is of the same complexity in both cases, but rather that if we can avoid checking all assignments in one case, we can avoid it in both. Nevertheless, the classical results on the P-completeness of these two problems are indeed the main ingredient we need, and our proof mostly consists of verifying that we can extend them to the new setting. We then go on to consider the prototypical W[P]-complete problem of the third statement. It is natural to expect that since the first and third statement refer to circuit satisfiability problems, without any particular restriction on the circuits, it should be possible to transform one to the other, and indeed we are able to do so using standard techniques, similar to what we did in Theorem 4.1.
Proof 6.2 (Proof of Theorem 6.1).
The proof is given in the four lemmas below which show the implications $1\Rightarrow 2$ (Lemma 6.3), $2\Rightarrow 1$ (Lemma 6.5), $3\Rightarrow 1$ (Lemma 6.7), $1\Rightarrow 4$ (Lemma 6.9).
@@ 6.1 Circuit-SAT and Horn Backdoors
Statement 1 of Theorem 6.1 implies statement 2.
@@
Statement 2 of Theorem 6.1 implies statement 1.
@@
The high-level idea is now to revisit the standard proof that Horn-SAT is P-complete, which relies on a reduction from the Circuit Value Problem. More precisely, assume that the given circuit is represented by a DAG that contains only $\land$,$\lor$ , ∨, and $\neg$ gates of maximum fan-in $2$. We first transform the circuit (in polynomial time) so that $\neg$ gates only appear directly connected to the inputs. Since the circuit is a DAG, we can assume that the gates are topologically sorted. We construct $\phi$ as follows:
@@
    For each input, $\land$, and $\lor$ gate $c$ of $C$, we construct two variables: $c^{p}$ and $c^{n}$. Intuitively, in circuit assignments that set $c$ to True, $c^{p}$ must be set to True and $c^{n}$ to False, while the opposite must happen for circuit assignments that set $c$ to False. Thus, for each such gate we add the clause ($(\neg c^{p}\lor\neg c^{n})$.
@@
    For each conjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses ($(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p}),(\neg c_{1}^{n}\lor c^{n}),(\neg c_{2}^{n}\lor c^{n}),(\neg c^{p}\lor c_{1}^{p}),(\neg c^{p}\lor c_{2}^{p})$
@@
    For each disjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses ($(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n}),(\neg c_{1}^{p}\lor c^{p}),(\neg c_{2}^{p}\lor c^{p}),(\neg c^{n}\lor c_{1}^{n}),(\neg c^{n}\lor c_{2}^{n})$
@@
    $c$ is a conjunction that has value True in $C$. Then $c_{1},c_{2}$ must evaluate to True in $C$, since these gates satisfy the property we have $c_{1}^{p},c_{2}^{p}$ set to True, so the clause ($(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
@@
    $c$ is a conjunction that has value False in $C$. Then one of $c_{1},c_{2}$ must evaluate to False in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{n}$ set to True. So, the clause ($(\neg c_{1}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.
@@
    $c$ is a disjunction that has value True in $C$. Then one of $c_{1},c_{2}$ must evaluate to True in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{p}$ set to True. So, the clause ($(\neg c_{1}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
@@
    $c$ is a disjunction that has value False in $C$. Then $c_{1},c_{2}$ must evaluate to False in $C$, since these gates satisfy the property we have $c_{1}^{n},c_{2}^{n}$ set to True, so the clause ($(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.

### v2

@@ 5 Modulators to Logarithmic Pathwidth and 2-SAT Backdoors
    ($\log\!\textrm{pw}$M-SETH is false) There is an algorithm that takes as input a CNF formula $\phi$, a modulator $M$ of size $m$, and a path decomposition of the primal graph of $\phi-M$ of width $O(\log|\phi|)$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
@@
We are therefore in a situation where, if we attempt to go through reachability, one problem we are dealing with has a quantification of the form $\exists\exists$ (for logarithmic pathwidth modulators the question is if there exists an assignment to the modulator so that the resulting graph has an $s\to t$ path) and the other has the form $\exists\forall$ (does there exist an assignment to the backdoor so that in the graph representing the 2-SAT implications there is no cycle creating a contradiction). What we need is to change the inner quantifier without knowing the value of the variables of the first quantifier block. We are thus led to the following definition of two auxiliary problems:
@@
Our strategy is now the following: we will show that Ann-s-t-Reach is equivalent to SAT parameterized by a logarithmic pathwidth modulator (intuitively this is easy, because both problems have an $\exists\exists$ quantification); we will show that SAT parameterized by a 2-SAT backdoor is equivalent to Ann-s-t-Non-Reach (again, both problems have an $\exists\forall$ quantification). Finally, we will revisit the proof of the Immerman–Szelepcsényi theorem and show that the key arguments still go through. In particular, the Immerman–Szelepcsényi theorem can be seen as a reduction from reachability to its complement. What we need to show is that this reduction can still be executed if arcs are annotated, as in our problems, in a way that preserves the answer for all assignments to $M$.
Proof 5.3 (Proof of Theorem˜5.1).
We split the proof into three parts. In Lemma˜5.4 we show an equivalence between 2-SAT backdoors and Ann-s-t-Non-Reach; in Lemma˜5.10 we show an equivalence between SAT with a modulator to logarithmic pathwidth and Ann-s-t-Reach; and in Lemma˜5.12 we show that Ann-s-t-Reach and Ann-s-t-Non-Reach have the same complexity by revisiting the proof of the Immerman–Szelepcsényi theorem.
@@
Proof $5.$
The proof is broken down into Lemma˜5.6 and Lemma˜5.8.
@@
We use a standard reduction of non-reachability to 2-SAT, except we have to take into account the arc annotations. Suppose we are given a Ann-s-t-Non-Reach instance. Let $G=(V,E)$ be the given DAG on $n$ vertices. We construct a CNF formula with variables the variables of $M$, given with the Ann-s-t-Non-Reach instance, to which we add a variable $x_{v}$ for each $v\in V$. We will set $M$ to be the backdoor set. Now, for each arc $(v_{1},v_{2})\in E$ we add the clause $(\neg x_{v_{1}}\lor x_{v_{2}})$ to the formula if the arc is not annotated, otherwise we add the clause $(\neg x_{v_{1}}\lor x_{v_{2}}\lor\neg\ell)$ if the arc is annotated with the literal $\ell$. We also add the unit clauses $(x_{s})$ and $(\neg x_{t})$. This completes the construction.
We now claim that if the formula we constructed is satisfiable, using the same assignment to $M$ yields a digraph where there is no path from $s$ to $t$. Indeed, suppose that for the same assignment there is a path from $s$ to $t$, meaning that the assignment sets to True all the literals used as annotations to the edges of the path. This implies that in the corresponding clauses, the backdoor literals we have added are set to False by the satisfying assignment. But then this would create a chain of implications given $x_{s}\to x_{t}$, which in conjunction with the two unit clauses would falsify the formula.
@@
Suppose now that the original CNF is satisfiable. We want to show that by using the same assignment to $M$ we eliminate all $s\to t$ paths. Indeed, suppose we find such a path, which for some $i\in[n]$ must go through $D_{i,0},D_{i,1}$. In particular, this means we have a path from $x_{i}$ to $\neg x_{i}$ and a path from $\neg x_{i}$ to $x_{i}$ in $D$ while respecting the annotations. However, recall that for each clause we have added a path of annotated arcs containing the negations of all backdoor literals of the clause, meaning that if we can follow an arc, the assignment to $M$ must be setting all the backdoor literals to False. We conclude that the assignment cannot be satisfying to the formula, as we can infer the implications $(x_{i}\to\neg x_{i})$ and $(\neg x_{i}\to x_{i})$.
@@ 5.2 Log-Pathwidth and Reachability
There is an algorithm solving Ann-s-t-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$, for some $\varepsilon>0$ if and only if the $\log\!\textrm{pw}$M-SETH is false.
@@
Let us now move to the other half of the lemma. Suppose there is an algorithm deciding satisfiability with a modulator $M$ to pathwidth $O(\log|\phi|)$ and we want to use this algorithm to solve Ann-s-t-Reach. We construct a formula which retains the set $M$ of annotation variables as the modulator. Given a digraph $G=(V,E)$ with $n$ vertices we assume without loss of generality that $n$ is a power of $2$ (otherwise add some dummy vertices). We construct a formula with $n\log n$ extra variables, called $x_{i,j}$, $i\in[n],j\in[\log n]$. The intended meaning is that the $\log n$ variables $x_{i,1},\ldots,x_{i,\log n}$ encode the index of the $i$-th vertex we would visit in an $s\to t$ path. To enforce this meaning we assume the vertices of $G$ are numbered from 0 to n1n-1 and with each vertex we associate a distinct truth assignment to $\log n$ variables by reading these variables as a binary number. We add clauses enforcing the variables $x_{1,j},j\in[\log n]$ to take the value corresponding to $s$ and the variables $x_{n,j},j\in[\log n]$ to take the value corresponding to $t$. For each $i\in[n-1]$ we add at most $n^{2}$ clauses to the formula as follows: for each $(u,v)\not\in E$ such that $u\neq v$, we add a clause ensuring that if the $x_{i,j}$ variables encode $u$ and the $x_{i+1,j}$ variables encode $v$, then the formula is not satisfied; for each $(u,v)\in E$ which is annotated with a literal $\ell$ we add the same clause (which is falsified if $x_{i,j},x_{i+1,j}$ encode $u,v$), but we add to this clause the literal $\ell$. If $(u,v)\in E$ is not annotated or if $u=v$, we do not add a clause. It is not hard to see that, if we remove the variables of $M$, this formula has pathwidth $O(\log n)$, because we can make n1n-1 bags, each containing $x_{i,j},x_{i+1,j},j\in[\log n]$ and this covers all edges.
To argue for correctness, if there is an $s\to t$ path in the original instance for some assignment to $M$ we use the same assignment in $\phi$ and extend it to a satisfying assignment as follows: if the path visits $r\leq n$ vertices in total, we set for each $i\in[r]$ the assignment to $x_{i,j},j\in[\log n]$ to encode the $i$-th vertex of the path; all remaining groups encode $t$. This satisfies the formula. For the converse direction, we can infer from a satisfying assignment to the formula a path of length at most n1n-1 by interpreting the assignment to $x_{i,j},j\in[\log n]$ as the encoding of the $i$-th vertex of the path, and the clauses we have added ensure that the path is valid, that is we are only followed to stay in the same vertex, follow an arc without annotation, or follow an arc whose annotation is set to True by the assignment to $M$.
@@ 5.3 Reachability and Non-Reachability
Before we begin, let us slightly edit $G_{1}$. Take $n=|V_{1}|$ copies of $V_{1}$, constructing for each $v\in V_{1}$ the vertices $v^{1},\ldots,v^{n}$. For each $i\in[n-1]$ and $v\in V_{1}$ we add the arc $(v^{i},v^{i+1})$ without annotation; for each arc $(u,v)$ and each $i\in[n-1]$ we add the arc $(u^{i},v^{i+1})$, maintaining the annotation of $(u,v)$. Finally, for the vertex $s^{n-1}$ and all $v\in V_{i}\setminus\{t\}$ we add the non-annotated arc $(s^{n-1},v^{n})$. This transformation has achieved the following: for each assignment to $M$, an $s\to t$ path exists in $G_{1}$ if and only if an $s^{1}\to t^{n}$ path exists in the new graph; there is a path from $s_{1}$ to $v^{n}$ for all $v\in V_{1}\setminus\{t\}$. As a result the question of whether an $s\to t$ path exists in $G_{1}$ is equivalent to asking whether the number of vertices reachable from $s^{1}$ in $\{v^{n}\ |\ v\in V_{1}\}$ is equal to $n$. Call the new graph $G_{1}^{\prime}$ and we want to construct $G_{2}$ such that for all assignments to $M$, there is a path $s_{2}\to t_{2}$ in $G_{2}$ if and only if n1n-1 vertices of $\{v^{n}\ |\ v\in V_{1}\}$ are reachable from $s^{1}$ in $G_{1}^{\prime}$. In the remainder, when we refer to the $i$-th layer of $G_{1}^{\prime}$ we will mean the set $\{v^{i}\ |v\in V_{1}\}$. Observe that $G_{1}^{\prime}$ has the property that all arcs go from a layer $i$ to a layer $i+1$.
We begin the construction of $G_{2}$ with $n^{2}$ vertices $s_{i,j}$ for $i,j\in[n]$, plus a source $s_{2}$ and a sink $t_{2}$. The intended meaning is that (for all assignments to $M$) $s_{i,j}$ will be reachable from $s_{2}$ if and only if there are exactly $j$ vertices of $G_{1}^{\prime}$ in layer $i$ which are reachable from $s^{1}$. We add a non-annotated arc $(s_{n,n-1},t_{2})$ and a non-annotated arc $(s_{2},s_{1,1})$. For each $i,j\in[n]$ we want to construct a checker gadget $C_{i,j}$ such that there will be an arc from $s_{i,j}$ to $C_{i,j}$ and arcs from $C_{i,j}$ to $s_{i+1,j^{\prime}}$ for some $j^{\prime}\in[n]$. We want to achieve the following property for all assignments $\sigma$ to $M$. Suppose that under some assignment $\sigma$, there are exactly $j$ vertices of layer $i$ of $G_{1}^{\prime}$ reachable from $s^{1}$. Then, there will be a path from $s_{i,j}$ to $s_{i+1,j^{\prime}}$ through $C_{i,j}$ if and only if the number of vertices of layer $i+1$ of $G_{1}^{\prime}$ reachable from $s^{1}$ is $j^{\prime}$. If we achieve this property we are done, as any path starts from $s_{1,1}$ and it is the case that exactly $1$ vertex of layer $1$ is reachable from $s^{1}$ ($s^{1}$ itself), so we can proceed by induction and we see that there is a path to $s_{n,n-1}$ (hence to $t_{2}$) if and only if n1n-1 vertices of the last layer of $G_{1}^{\prime}$ are reachable. Hence, there will be a $s_{2}\to t_{2}$ path in $G_{2}$ if and only if there is no $s_{1}\to t_{1}$ path in $G_{1}$.
@@ 6 Circuit-SAT and Horn Backdoors
Before we go on let us give some intuition. Similarly to Section˜5, it will be helpful to recall some basic facts from computational complexity theory. For the first two statements, we are dealing with the satisfiability problem for circuits and SAT with a Horn backdoor. Intuitively, it is natural to expect these two problems to be equivalent because once we fix either the input to a given circuit or the assignment to the Horn backdoor, deciding if the circuit is satisfied or if the backdoor assignment can be extended to a satisfying assignment are problems with the same classical complexity, that is, they are P-complete [44]. As in Section˜5 this intuition is not in itself sufficient to imply the equivalence, because what we want is not to claim that once we have an assignment, checking it is of the same complexity in both cases, but rather that if we can avoid checking all assignments in one case, we can avoid it in both. Nevertheless, the classical results on the P-completeness of these two problems are indeed the main ingredient we need, and our proof mostly consists of verifying that we can extend them to the new setting. We then go on to consider the prototypical W[P]-complete problem of the third statement. It is natural to expect that since the first and third statement refer to circuit satisfiability problems, without any particular restriction on the circuits, it should be possible to transform one to the other, and indeed we are able to do so using standard techniques, similar to what we did in Theorem˜4.1.
Proof 6.2 (Proof of Theorem˜6.1).
The proof is given in the four lemmas below which show the implications $1\Rightarrow 2$ (Lemma˜6.3), $2\Rightarrow 1$ (Lemma˜6.5), $3\Rightarrow 1$ (Lemma˜6.7), $1\Rightarrow 4$ (Lemma˜6.9).
@@ 6.1 Circuit-SAT and Horn Backdoors
Statement 1 of Theorem˜6.1 implies statement 2.
@@
Statement 2 of Theorem˜6.1 implies statement 1.
@@
The high-level idea is now to revisit the standard proof that Horn-SAT is P-complete, which relies on a reduction from the Circuit Value Problem. More precisely, assume that the given circuit is represented by a DAG that contains only $\land,\lor$, and $\neg$ gates of maximum fan-in $2$. We first transform the circuit (in polynomial time) so that $\neg$ gates only appear directly connected to the inputs. Since the circuit is a DAG, we can assume that the gates are topologically sorted. We construct $\phi$ as follows:
@@
    For each input, $\land$, and $\lor$ gate $c$ of $C$, we construct two variables: $c^{p}$ and $c^{n}$. Intuitively, in circuit assignments that set $c$ to True, $c^{p}$ must be set to True and $c^{n}$ to False, while the opposite must happen for circuit assignments that set $c$ to False. Thus, for each such gate we add the clause $(\neg c^{p}\lor\neg c^{n})$.
@@
    For each conjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses $(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p}),(\neg c_{1}^{n}\lor c^{n}),(\neg c_{2}^{n}\lor c^{n}),(\neg c^{p}\lor c_{1}^{p}),(\neg c^{p}\lor c_{2}^{p})$
@@
    For each disjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses $(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n}),(\neg c_{1}^{p}\lor c^{p}),(\neg c_{2}^{p}\lor c^{p}),(\neg c^{n}\lor c_{1}^{n}),(\neg c^{n}\lor c_{2}^{n})$
@@
    $c$ is a conjunction that has value True in $C$. Then $c_{1},c_{2}$ must evaluate to True in $C$, since these gates satisfy the property we have $c_{1}^{p},c_{2}^{p}$ set to True, so the clause $(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
@@
    $c$ is a conjunction that has value False in $C$. Then one of $c_{1},c_{2}$ must evaluate to False in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{n}$ set to True. So, the clause $(\neg c_{1}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.
@@
    $c$ is a disjunction that has value True in $C$. Then one of $c_{1},c_{2}$ must evaluate to True in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{p}$ set to True. So, the clause $(\neg c_{1}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
@@
    $c$ is a disjunction that has value False in $C$. Then $c_{1},c_{2}$ must evaluate to False in $C$, since these gates satisfy the property we have $c_{1}^{n},c_{2}^{n}$ set to True, so the clause $(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.